# XAI4LLM. Let Machine Learning Models and LLMs Collaborate for Enhanced In-Context Learning in Healthcare

### v2

@@ 1. Introduction
Context. Unlike traditional machine-learning models, which are often limited to functioning as “narrow-expert” systems optimized for specific tasks, modern generative models demonstrate a remarkable ability to learn from and sample a broad range of data distributions, for example, those obtained from vast, diverse sources from large-scale internet, encompassing text and visual content. This capability enables them to generate authentic-looking data that closely resembles the training data. Furthermore, recently, the advancement of natural language processing (NLP), combined with increased computational power and data availability, has spurred the development of large-scale generative models, known as large language models (LLMs), such as ChatGPT (OpenAI, 2022), and Gemini (Team et al., 2023). This new form of learning does not rely on limited, narrowly defined datasets to support a traditional $train\rightarrow prediction\rightarrow label$ paradigm (in a supervised setting). Instead, pretrained large language models (pLMS) leverage massive parameter data to facilitate a direct $prediction\rightarrow label$ process, potentially conserving resources and enhancing nuanced learning. Consequently, once provided with the appropriate context and domain knowledge, these models can be optimized for use in various downstream tasks. Overall, thanks to their wide-ranging, generalist emergent capabilities (Schaeffer et al., 2023), these models have introduced a myriad of opportunities for research across various AI disciplines (Imani et al., 2023; Zhang et al., 2023; Jiang et al., 2023; Xu et al., 2023; Li et al., 2023).
@@
Communication Styles with LLMs. We recognize in this work, the impact of multi-turn (conversational) versus single-turn interactions on LLMs’ diagnostic precision. Communication styles with LLMs leverage structured interaction patterns, employing “<User>” and “<System>” tags to facilitate conversations between the user and the LLM. Two predominant styles are recognized: multi-turn (“conversational”) and single-turn interactions. The conversational style utilizes a sequence of short, interactive exchanges leading to a heart disease risk assessment, enabling incremental information gathering similar to the approach of a clinician approach with a patient, while the latter compiles all patient details into one (or few) comprehensive prompts. The challenge revolves around determining which communication style — the dynamic, detail-oriented conversational or the direct, comprehensive single-turn — effectively enhances diagnostic precision of the LLMs.
@@ 1.1. Research Objectives and Contributions
    Optimizing Communication Styles and Profiles in LLM-based Diagnostics. Our research recognizes two communication styles, intricately combined with patient profile construction: (i) the “conversational” style, which employs multi-turn interactions utilizing numerical data to mimic the dynamic, incremental nature of clinician-patient dialogues, and (ii) the “Single-Long Prompt” style, which integrates all patient details into a comprehensive narrative for a holistic assessment similar to a review of clinicians of a complete medical file before diagnosis. This approach is motivated by the need to understand how the structured presentation of patient data—whether through a detailed, step-by-step conversation or a broad, single overview—can effectively harness the capabilities of LLM to yield accurate heart disease risk assessments. Examples of these two approaches are shown in the following section.
@@ 3.2.1. Discriminative approach to binary classification
Discriminative models, or direct learning models, seek to map input features $X$ to labels $Y$ via a function $f$:
(4) $f:X\rightarrow Y$
In this approach, the objective of the models –often referred to as a learning model– is to learn the mapping $f$ to accurately predict the label of unseen data points. This is achieved by minimizing a loss function, such as cross-entropy, which quantifies the discrepancy between the predicted labels and the actual labels.
@@ 5.1. Key Evaluation Metrics
    For General Performance Evaluation: We use $\beta=1$ (yielding the F1 score) to balance precision and recall. This choice reflects our interest in evaluating the performance of the model where both false positives and false negatives carry similar weight.
    (7) $F_{1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
@@
    For Risk Sensitivity: Considering the critical nature of clinical outcomes, we also use $\beta=3$, yielding the F3 score. This score gives more weight to recall, reflecting the higher cost of false negatives in clinical settings.
    (8) $F_{3}=\frac{(1+3^{2})\times\text{Precision}\times\text{Recall}}{3^{2}\times\text{Precision}+\text{Recall}}$
By adopting the F3 score, we emphasize the importance of minimizing false negatives due to the grave consequences of overlooking at-risk individuals. This approach ensures that our model evaluation aligns with the prioritization of patient safety and effective risk identification.
@@
In assessing fairness, we focus on the following metrics which are crucial for ensuring equitable treatment across different demographic groups.
Equality of Opportunity ($\Delta_{EO}$) Equality of opportunity is assessed by the difference in true positive rates between groups $\Delta_{EO}=|\text{TPR}_{M}-\text{TPR}_{F}|$ with TPR representing the true positive rate. Predictive Parity ($\Delta_{PP}$) Predictive parity requires that different groups have equal positive predictive values: $\Delta_{PP}=|\text{PPV}_{M}-\text{PPV}_{F}|$, where PPV is the positive predictive value. Predictive Equality Accuracy ($\Delta_{PEAc}$). Finally, the difference in accuracy across genders is defined by $\Delta_{PEAc}=|\text{Acc}_{M}-\text{Acc}_{F}|$.
@@ 6. Results
Domain Knowledge NC Style NL-St Style
Model XAI-model Acc. Pre. Rec. $F_{\beta}$ Acc. Pre. Rec. $F_{\beta}$
@@
Majority Zeros 0.402 0.300 0.059 0.098 0.064 0.402 0.300 0.059 0.098 0.064
Majority Ones 0.489 0.524 0.843 0.647 0.795 0.489 0.524 0.843 0.647 0.795
@@ 7. Conclusion
This research focuses on integrating large language models (LLMs) into healthcare decision-making, especially for heart risk prediction through zero/few-shot in-context learning (ICL). The primary contribution of this work is the development of a systematic structured prompt construction that has been shown enhancing diagnostic accuracy of LLMs by incorporating domain-specific medical knowledge into their decision-making process. This includes creating multi-layered structured templates that efficiently integrate numerical and natural language user profiles along with critical medical domain knowledge. Furthermore, this study enhances our understanding of the optimal communication styles with LLMs—whether conversational or single-long prompt—that can best utilize LLMs in clinical diagnostics.
To provide a clearer summary of the results and key takeaways, we present the main findings below and in Figure 5, which includes a checklist of the primary results obtained.
@@
The results of the study are promising, demonstrating that the structured approach to prompt construction and the nuanced integration of communication styles can significantly impact the diagnostic precision of LLMs in healthcare. By comparing these new ways of prediction compared to traditional ML models, the research highlights the potential of LLMs to revolutionize medical diagnostics by offering more adaptive, nuanced, and potentially resource-conserving methods. Additionally, it emphasizes the importance of addressing inherent risks such as false negatives and biases, which are critical for ensuring the safe and equitable application of LLMs in clinical settings. Through these insights and methodologies, this research paves the way for more effective and nuanced applications of AI in healthcare.
@@
Summary of LLM Diagnostic Performance Analysis Main Findings: The graph in Figure 5 illustrates the relationship between the length of communication or prompt (synonymous with the number of few-shot examples, ranging from 0 to 32) and the resulting diagnostic performance measured by an accuracy/cost-sensitive accuracy metric. (1) Conversational Style Achieves Higher Performance. The “Numerical Conversational (NC)” style shows improvement with an increasing number of few-shot examples and surpassed the “Natural language Single Turn (NL-St)” prompt style in achieving peak accuracy, indicating that conversational interaction is potentially more effective for long LLM diagnostic description. (2) CoT Reasoning Aids Textual Presentation. Chain-of-thought reasoning maintains quality performance with growing example counts, within the narrative single-prompt context. This advantage points to effectiveness of COT in helping LLMs manage and utilize larger amounts of textual data without a decline in diagnostic accuracy. Surprisingly, COT has an adverse impact in the conversational mode. (3) The Adverse Impact of Long Prompts on Quality of Predictions. In contrast to the conversational model, which appears less sensitive to the number of examples, the quality of the NL-St profile declines with increasing prompt length. This suggests that the conversational mode is better suited to digest a larger volume of examples, processing them sequentially, one by one, and may imply that it is better equipped to manage and interpret extensive data sets.

### v3

@@ 1. Introduction
Context. Unlike traditional machine-learning models, which are often limited to functioning as “narrow-expert” systems optimized for specific tasks, modern generative models demonstrate a remarkable ability to learn from and sample a broad range of data distributions, for example, those obtained from vast, diverse sources from large-scale internet, encompassing text and visual content. This capability enables them to generate authentic-looking data that closely resembles the training data. Furthermore, recently, the advancement of natural language processing (NLP), combined with increased computational power and data availability, has spurred the development of large-scale generative models, known as large language models (LLMs), such as ChatGPT (OpenAI, 2022), and Gemini (Team et al., 2023; Li et al., 2023). This new form of learning does not rely on limited, narrowly defined datasets to support a traditional $train\rightarrow prediction\rightarrow label$ paradigm (in a supervised setting). Instead, pretrained large language models (pLMS) leverage massive parameter data to facilitate a direct $prediction\rightarrow label$ process, potentially conserving resources and enhancing nuanced learning. Consequently, once provided with the appropriate context and domain knowledge, these models can be optimized for use in various downstream tasks. Overall, thanks to their wide-ranging, generalist emergent capabilities (Schaeffer et al., 2023), these models have introduced a myriad of opportunities for research across various AI disciplines (Imani et al., 2023; Zhang et al., 2023; Jiang et al., 2023; Xu et al., 2023; Sharma et al., 2024; Chang et al., 2024).
@@
Communication Styles with LLMs. We recognize in this work, the impact of multi-turn (conversational) versus single-turn interactions on LLMs’ diagnostic precision. Communication styles with LLMs leverage structured interaction patterns, employing “¡User¿” and “¡System¿” tags to facilitate conversations between the user and the LLM. Two predominant styles are recognized: multi-turn (“conversational”) and single-turn interactions. The conversational style utilizes a sequence of short, interactive exchanges leading to a heart disease risk assessment, enabling incremental information gathering similar to the approach of a clinician approach with a patient, while the latter compiles all patient details into one (or few) comprehensive prompts. The challenge revolves around determining which communication style — the dynamic, detail-oriented conversational or the direct, comprehensive single-turn — effectively enhances diagnostic precision of the LLMs.
@@ 1.1. Research Objectives and Contributions
    Optimizing Communication Styles and Profiles in LLM-based Diagnostics. Our research recognizes two communication styles, intricately combined with patient profile construction: (i) the “conversational” style, which employs multi-turn interactions utilizing numerical data to mimic the dynamic, incremental nature of clinician-patient dialogues, and (ii) the “Single-Long Prompt” style, which integrates all patient details into a comprehensive narrative for a holistic assessment similar to a review of clinicians of a complete medical file before diagnosis. This approach is motivated by the need to understand how the structured presentation of patient data—whether through a detailed, step-by-step conversation or a broad, single overview—can effectively harness the capabilities of LLM to yield accurate heart disease risk assessments. Examples of these two approaches are shown below.
@@ 3.2.1. Discriminative approach to binary classification
Discriminative models, or direct learning models, seek to map input features $X$ to labels $Y$ via a function $f:X\rightarrow Y$. In this approach, the objective of the models –often referred to as a learning model– is to learn the mapping $f$ to accurately predict the label of unseen data points. This is achieved by minimizing a loss function, such as cross-entropy, which quantifies the discrepancy between the predicted labels and the actual labels.
@@ 5.1. Key Evaluation Metrics
    For General Performance Evaluation: We use $\beta=1$ (yielding the F1 score) to balance precision and recall. This choice reflects our interest in evaluating the performance of the model where both false positives and false negatives carry similar weight $F_{1}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$
@@
    For Risk Sensitivity: Considering the critical nature of clinical outcomes, we also use $\beta=3$, yielding the F3 score. This score gives more weight to recall, reflecting the higher cost of false negatives in clinical settings $F_{3}=\frac{(1+3^{2})\times\text{Precision}\times\text{Recall}}{3^{2}\times\text{Precision}+\text{Recall}}$. By adopting the F3 score, we emphasize the importance of minimizing false negatives due to the grave consequences of overlooking at-risk individuals. This approach ensures that our model evaluation aligns with the prioritization of patient safety and effective risk identification.
@@
In assessing fairness, we focus on the following metrics which are crucial for ensuring equitable treatment across different demographic groups. Equality of Opportunity ($\Delta_{EO}$) Equality of opportunity is assessed by the difference in true positive rates between groups $\Delta_{EO}=|\text{TPR}_{M}-\text{TPR}_{F}|$ with TPR representing the true positive rate. Predictive Parity ($\Delta_{PP}$) Predictive parity requires that different groups have equal positive predictive values: $\Delta_{PP}=|\text{PPV}_{M}-\text{PPV}_{F}|$, where PPV is the positive predictive value. Predictive Equality Accuracy ($\Delta_{PEAc}$). Finally, the difference in accuracy across genders is defined by $\Delta_{PEAc}=|\text{Acc}_{M}-\text{Acc}_{F}|$.
@@ 6. Results
DK. NC Style NL-St Style
Model XAI Acc. Pre. Rec. $F_{\beta}$ Acc. Pre. Rec. $F_{\beta}$
@@
Maj.Zeros 0.402 0.300 0.059 0.098 0.064 0.402 0.300 0.059 0.098 0.064
Maj.Ones 0.489 0.524 0.843 0.647 0.795 0.489 0.524 0.843 0.647 0.795
@@ 7. Conclusion
This research integrates large language models (LLMs) into healthcare decision-making, particularly for heart risk prediction using zero/few-shot in-context learning (ICL). The main contribution is the development of a systematic structured prompt construction that enhances diagnostic accuracy in ICL by incorporating domain-specific medical knowledge. This involves creating multi-layered templates that combine numerical and natural language user profiles with critical medical knowledge. The study also explores the best communication styles with LLMs—conversational or single-long prompt—for clinical diagnostics.
@@
The results are promising, showing that structured prompt construction and nuanced communication styles can significantly improve diagnostic precision of LLMs in healthcare. By comparing these methods to traditional ML models, the research demonstrates LLMs’ potential to impact medical diagnostics with more adaptive and resource-efficient methods. It also stresses the importance of addressing risks such as false negatives and biases to ensure safe and equitable use of LLMs in clinical settings.
@@
Summary of LLM Diagnostic Performance Analysis Main Findings: The graph in Figure 5 illustrates the relationship between the length of communication or prompt (synonymous with the number of few-shot examples, ranging from 0 to 32) and the resulting diagnostic performance measured by an accuracy/cost-sensitive accuracy metric. Conversational Style Achieves Higher Performance: The “Numerical Conversational (NC)” style shows improvement with an increasing number of few-shot examples and surpassed the “Natural language Single Turn (NL-St)” prompt style in achieving peak accuracy, indicating that conversational interaction is potentially more effective for long LLM diagnostic descriptions. CoT Reasoning Aids Textual Presentation: Chain-of-thought reasoning maintains quality performance with growing example counts within the narrative single-prompt context. This advantage points to the effectiveness of CoT in helping LLMs manage and utilize larger amounts of textual data without a decline in diagnostic accuracy. Surprisingly, CoT has an adverse impact in the conversational mode. The Adverse Impact of Long Prompts on Quality of Predictions: In contrast to the conversational model, which appears less sensitive to the number of examples, the quality of the NL-St profile declines with increasing prompt length. This suggests that the conversational mode is better suited to digest a larger volume of examples, processing them sequentially, one by one, and may imply that it is better equipped to manage and interpret extensive data sets.